OpenStudy (anonymous):

fx = π^6(8x − 1). f ' x = ?

OpenStudy (anonymous):

\[f(x)=\pi ^6(8x-1)^1\] \[f'(x)=1*\pi ^6 * 8=8\pi^6\] Take the derivative of the overall expression, then multiply by the derivative of what is in the parentheses

OpenStudy (anonymous):

\[\pi^6\] is simply a coefficient

OpenStudy (anonymous):

gotcha thanks